OpenStudy (anonymous):

The value of 5/(9x^2 + 6x + 11) is maximum when x=?

OpenStudy (anonymous):

9 x^2 + 6x + 11 should be minimum

OpenStudy (anonymous):

take derivative

OpenStudy (anonymous):

I dont think so

OpenStudy (anonymous):

You should take derivative first..

OpenStudy (anonymous):

or complete the squre for 9 x^2 + 6x + 11

OpenStudy (anonymous):

18 x + 6

OpenStudy (anonymous):

well so x=-2 will minimize 9 x^2 + 6x + 11 and consequencly will maximize 5/9 x^2 + 6x + 11

OpenStudy (anonymous):

hw u got x = -2

OpenStudy (anonymous):

The answer should be x = -1/3

OpenStudy (anonymous):

sorry 18 x + 6=0 then x=-1/3

OpenStudy (anonymous):

got it thxxx

OpenStudy (anonymous):

Why was I sent this in a message? xD

OpenStudy (anonymous):

A plot of 5/(9x^2 + 6x + 11) and it's derivative, in blue and red respectively, is attached.